OpenStudy (anonymous):

one of the roots of the equation x2 + kx - 6 =0 is 3 , and k is constant.? which of the quantity is larger ? A. value of k B. -1

OpenStudy (anonymous):

plug in x=3: 9 - 3k - 6 = 0 3k = 3 k = 1 A

myininaya (myininaya):

9+3k-6=0 3+3k=0 k=-1

OpenStudy (anonymous):

k =-1 is correct

myininaya (myininaya):

A=B

OpenStudy (anonymous):

yup